# Multiclass Portfolio Optimization via Variational Quantum Eigensolver with Dicke State Ansatz

### v2

@@ III CMA-ES $\bm{70}$ $\mathbb{1}$ $335\pm 12$
Table 3: A summary of experiments results for each scenario considering only Dicke state ansatz. The column All represents the frequency of the optimal solution for each scenario considering $100$ experiments for each optimizer. $p(x^{*})\geq 0.95$ filter only the quantum outputs which the probability of the optimal state is greater than $0.95$. As can be seen, by applying this filter the frequency in CMAES drastically reduces in some scenarios (numbers indicated in red). The last column represents the average running time and the standard deviation of each optimizer.
Although the optimal solution frequently appears among the $100$ trials, this alone is insufficient to assess the quality of the quantum output. For that, we examine the probability of observing the target state, $p(x^{*})$ in each trial. As shown in Fig.5, depending on the optimizer and the scenario, $p(x^{*})$ often deviates significantly from $1$. For example, with the CMA-ES optimizer, the distribution of the probability of measuring the target state differs notably between scenarios: in scenario III, $p(x^{*})$ is broadly distributed with few cases near $1$, while in scenario II, $p(x^{*})\geq 0.95$ in almost all trials. This effect is further highlighted when filtering trials where $p(x^{*})\geq 0.95$, which substantially reduces the number of successful cases, except in scenario II, suggesting a less complex landscape (see Table 3).
It is clear that the optimizers guide the distribution to the right direction, this fact can be seen in Fig.5, where we compare the approximation ratio distribution before and after the parameter optimization process. Note that the values of the initial distribution are below $a_{r}=0.9$, but after optimization we can see a displacement to the right, which means that optimization succeeded in obtaining a set of parameters which generates a quantum state whose state distribution has a high probability of measuring a state with a high approximation ratio.
Note that even when the probability of measuring the target state is low, $a_{r}\gtrsim 0.9$ for most optimizers, as illustrated in the plot of Fig.5 (colored bars). This phenomenon is attributed to Eq.(7), which, in the quantum approach, serves as a weighted average that encapsulates the expected value of the Hamiltonian. This means that the approximation ratio of the quantum output distribution is a weighted sum of the approximation ratios of each state. Consequently, as the system approaches the optimal region after optimization, where the majority of states within the distribution have lower energy, the approximation ratio tends to be high.
@@ VI Conclusion
In this work, we introduce an innovative approach to exploring multiclass portfolio optimization through a parameterized version of multiple Dicke states within a Variational Quantum Eigensolver (VQE) framework. We analyze three distinct scenarios by varying the complexity of the search space and the number of parameters in the parameterized circuit. Our comparative evaluation indicates that Dicke states outperform other ansatzes, making them a suitable choice for this type of problem.
@@
Another intriguing aspect we began to evaluate is the relationship between state fidelity and optimal parameters, as discussed in the Appendix XIV of this paper. In future research, we aim to address the open questions highlighted here and further investigate the use of the Quantum Approximate Optimization Algorithm (QAOA) with Dicke states for multiclass portfolio optimization. This will include exploring different mixer Hamiltonians.
All our results were obtained through simulations, as our primary objective was to gain a deeper understanding of the algorithm’s potential for solving realistic problems. A future avenue of research will involve testing these findings on current quantum hardware, comparing the effectiveness of Zero-Noise Extrapolation (ZNE), Probabilistic Error Cancellation (PEC), and Conditional Value at Risk (CVaR) in mitigating errors during the computation of expectation values.
Despite these promising results, substantial research is still needed to fully understand the real impact of variational quantum algorithms on commercially significant problems. Our goal was not to provide definitive proof of this impact, but rather to test the viability of utilizing VQE in a more realistic financial scenario. We believe that our findings represent a meaningful step towards this objective.
Acknowledgements. We acknowledge the Instituto de Ciência e Tecnologia Itaú for all the support during this research project.
@@ VII Data Availability
The datasets generated and/or analysed during the current study are available in the Yahoo Finance and can be downloaded using the following python library: yfinance. The period of the datasets considered in this work are: 2023-12-18 until 2024-12-16 and 2024-01-02 until 2025-01-02.

### v3

@@ III CMA-ES $\bm{70}$ $\mathbb{1}$ $335\pm 12$
Table 3: A summary of experiments results for each scenario considering only Dicke state ansatz. The column All represents the frequency of the optimal global solution $(x^{*})$, among $100$ experiments, counted if it appears with highest probability. $p(x^{*})\geq 0.95$ filter only the quantum outputs which the probability of the optimal state is greater than $0.95$. As can be seen, by applying this filter the frequency in CMAES drastically reduces in some scenarios (numbers indicated in red). The last column represents the average running time and the standard deviation of each optimizer.
It is clear that the optimizers guide the distribution toward the right direction, this fact can be seen in Fig.5, where we compare the approximation ratio distribution before and after the parameter optimization process. Note that the values of the initial distribution are below $a_{r}=0.9$, but after optimization we can see a displacement to the right, which means that optimization succeeded in obtaining a set of parameters which generates a quantum state whose state distribution has a high probability of measuring a state with a high approximation ratio.
Note that even when the probability of measuring the target state is low (see probability of state of measure 0 in Fig. 6), $a_{r}\gtrsim 0.9$ for most optimizers, see Fig. 5 (colored bars). This phenomenon is linked to Eq.(7), which, from a quantum perspective, acts as a weighted average representing the expected value of the Hamiltonian. This implies that the approximation ratio of the quantum output distribution is essentially a weighted sum of the approximation ratios of individual states. As a result, when the system moves toward the optimal region post-optimization, where most states in the distribution have lower energy, the approximation ratio generally becomes higher.
@@
It is important to mention that recent studies [Scriva2024Challenges, ZhouLeo2020Quantum, maurizio2025genomics] suggest the potential quantum advantage of variational algorithms like VQE and QAOA arises when the number of function evaluations $(n_{calls}\equiv n_{shots}\times n_{iterations})$ remains smaller than the size of the effective search space ($n_{\text{search space}}$). By applying the Dicke state ansatz in our problem, we significantly reduce the search space from $2^{n}$ by $\prod_{i}^{n_{c}}C_{n_{i},k_{i}}\sim O(n^{k})$. However, for the set of experiments provided here, $n_{calls}\sim 4.096.000\gg n_{\text{search space}}$ as can be seen in the last column of Table 2.
@@ VI Conclusion
In this work, we introduce an innovative approach to exploring multiclass portfolio optimization through a parameterized version of multiple Dicke states within a VQE framework. We analyze three distinct scenarios by varying the complexity of the search space and the number of parameters in the parameterized circuit. Our comparative evaluation indicates that Dicke states outperform other ansatzes, making them a suitable choice for this type of problem.
@@
Another intriguing aspect we began to evaluate is the relationship between state fidelity and optimal parameters, as discussed in the Supplementary Material. In future research, our aim is to address the open questions highlighted here and to further investigate the use of the QAOA with Dicke states for multiclass portfolio optimization. This will include exploring different mixer Hamiltonians.
All our results were obtained through simulations, as our primary objective was to gain a deeper understanding of the algorithm’s potential for solving realistic portfolio optimization problems, encoding the diversification constraints in the quantum state preparation.
Although the proposed Dicke-state framework achieved promising results, further investigation is needed to evaluate its scalability and practical benefits in large-scale, real-world portfolio optimization. The potential advantage of variational algorithms emerges when the number of function evaluations remains smaller than the effective search space, motivating future studies combining the Dicke ansatz with CMA-ES and Conditional Value at Risk (CVaR) to explore possible regimes of quantum advantage. While current simulations assume all-to-all connectivity, the approach is hardware-agnostic and compatible with realistic 22D or sparse architectures, offering structural efficiency by reducing the feasible subspace and eliminating penalty terms. Preliminary hardware tests on IBM quantum devices can be seen in the Supplementary Material and confirm the expected sensitivity to noise and circuit depth, reinforcing the focus on noiseless simulations to isolate algorithmic behavior. Finally, quantum-optimized portfolios showed consistency with classical allocations in risk–return balance and diversification quality, encouraging future work on explicit financial performance metrics.
Future research is also needed to test these findings on current quantum hardware, comparing the effectiveness of Zero-Noise Extrapolation (ZNE), Probabilistic Error cancellation (PEC) and CVaR in mitigating errors during the calculation of expectation values. Our goal was not to provide definitive proof of this impact, but rather to test the viability of utilizing VQE in a more realistic financial scenario and open a new direction of exploration that is encoding optimization constraints in state preparation. We believe that our findings represent a significant step towards this objective.
Author contribution. SB and JV conceived the project. SB, JV and GM developed the theoretical framework. JV, GM and VB performed the experiment and collected the data. JV and SB analyzed the experimental data and prepared the figures. All authors contributed to the writing of the manuscript.
@@
Funding declaration. This research did not receive funding.
@@ VII Data Availability
The datasets generated and/or analyzed during the current study are available in the Yahoo Finance and can be downloaded using the following Python library: yfinance. The period of the datasets considered in this work are: 2023-12-18 until 2024-12-16 and 2024-01-02 until 2025-01-02.
@@ VIII Code Availability
Supplementary Methods
IX Experiment settings
The VQE algorithm simulations were performed using the instance ml.g5.xlarge on Amazon SageMaker. This machine has four virtual CPUs, 16 GB RAM and NVIDIA GPU A10G with 24 GB GDDR6 VRAM. This system runs on Linux, with Python version 3.9.5, CUDA version 12.4, NVIDIA Graphics driver version 550.144.03. The Python libraries used to perform the simulations are: qiskit version 1.3.2 [javadi2024quantum], qiskit-aer-gpu version 0.15.1, qiskit-algorithms version 0.3.1, qiskit-optimization version 0.6.1, numpy version 2.0.2, optuna version 4.1.0 and CMA-ES version 0.11.1. All initial points were generated with NumPy random methods using the seed $42$, with the exception for CMA-ES and Random Sampler due to optuna’s parameter sampling method. However, we guarantee that all parameters are in the interval $0,2\pi]$, independently of the method used to sample the parameters values. CMA-ES and Random Sampler were used through the Optuna interface, whereas COBYLA, SPSA and QNSPSA were used through the Qiskit algorithms package.
As mentioned in the main paper, we solved each scenario proposed with the classical optimizer SCIP and considered its results as a reference. These results are summarized in Table 4.
Scenario Objective function value Running time (s)
I -0.818106 0.01 $\pm$ 0.001
II -1.474237 0.03 $\pm$ 0.001
III -2.00332 0.03 $\pm$ 0.001
Table 4: A summary of the reference values computed using SCIP optimizer for each scenario. The column Objective function value holds the optimal values for each case. The column Running time represents the mean time spent and the standard deviation of 100 runs of the optimization problem.
X TwoLocal ansatz
The TwoLocal circuit is an ansatz that alternates rotation and entanglement layers. The rotation gates are applied in each qubit individually and the gates $RX$, RYRY, RZRZ, $U_{1}$, $U_{2}$ and $U_{3}$ are some examples of rotation gates that we can use in this ansatz. Meanwhile, the entanglement gates are used to create entanglement between pairs of qubits considered in each entanglement structure design, for instance: we can create entanglement between all qubits or just in some qubits in a way that connects the qubits through a line. In order to create the desired entanglement structure, we can use gates such as: $CX$, CYCY, CZCZ, CRXCRX, CRYCRY and $CRZ$. For instance, in Table 5 we show some TwoLocal settings used in Scenario I experiments.
Name Qubits Parameters Rotation Blocks Entanglement Blocks Entanglement Repetitions Skip Final Rotation Layer
TwoLocal1 10 30 3 x RY CX Full 1 True
TwoLocal2 10 30 1 x RY CX Full 2 True
TwoLocal3 10 30 1 x RY CX Full 3 False
TwoLocal4 10 30 3 x RY CX Linear 1 True
TwoLocal5 10 30 1 x RY CX Linear 2 True
TwoLocal6 10 30 1 x RY CX Linear 3 False
TwoLocal7 10 30 3 x RY CX Reverse Linear 1 True
TwoLocal8 10 30 1 x RY CX Reverse Linear 2 True
TwoLocal9 10 30 1 x RY CX Reverse Linear 3 False
TwoLocal10 10 30 3 x RY CX Circular 1 True
TwoLocal11 10 30 1 x RY CX Circular 2 True
TwoLocal12 10 30 1 x RY CX Circular 3 False
TwoLocal13 10 30 3 x RY CX SCA 1 True
TwoLocal14 10 30 1 x RY CX SCA 2 True
TwoLocal15 10 30 1 x RY CX SCA 3 False
TwoLocal16 10 30 3 x RY CX Pairwise 1 True
TwoLocal17 10 30 1 x RY CX Pairwise 2 True
TwoLocal18 10 30 1 x RY CX Pairwise 3 False
Table 5: Settings used to compose each TwoLocal variant consired in the experiments in Scenario I. All circuits for Scenario I were created to have 30 parameters and 10 qubits, then we just varied the amount of rotation blocks, entanglement structure, the number of layers repetitions and whether we skip the final rotation layer. In this case we only used RYRY and CXCX gates because relative phases are not relevant.
XI Derivation of the number of parameters in the parametrized Dicke State
The total number of parameters is directly related to the number of gates $CRY$ and $CCRY$, and each of these gates has one free parameter. We defined a range of values for $n$ and considered [1,n1][1,n-1] as the range for $k$. Then, we constructed the Dicke state ansatz for all values of $n$ and $k$ defined in the ranges mentioned above and computed the total number of parameterized gates. With these data in our hands, we were able to plot the curves (see Figure 8) and also perform a regression to derive a formula that dictates the relationship between the number of qubits and the Hamming weight, with the number of parameters in the Dicke state ansatz.
Taking into account a fixed value for $k$ and varying the number of qubits $n$ in the range $[k+1,d]$, where $d>k+1$, and building the Dicke state ansatz circuit for each value of $n$, we can calculate the number of parameters. Plotting $n_{p}$ by $n$, we saw a polynomial function of degree 1, then we can write the following expression
$n_{p}=mn+b.$ (8)
The angular coefficient can be calculated by the following equation
$m=\frac{n_{p}^{f}-n_{p}^{i}}{n_{f}-n_{i}},$ (9)
computing the value of $m$ with the data obtained from the quantum circuits of Dicke state ansatz, we found that $m=k$, which implies that the angular coefficient is equal to the Hamming weight of the Dicke state. The linear coefficient can be computed using the following expression
$b=n_{p}^{i}-kn_{i},$ (10)
and in order to find out if there is a general rule for $b$, we tested different values of $k$ for a fixed $d$ and we found that the value of the linear coefficient is given by
$b=\frac{k(k+1)}{2}.$ (11)
Thus, combining all the elements computed above, we obtain the following equation for a Dicke state ansatz with $n$ qubits and Hamming weight $k$
$n_{p}=kn-\frac{k(k+1)}{2}.$ (12)
Figure 8: Scaling of the number of parameters in Dicke state ansatz for different numbers of qubits and varying the Hamming weight $k$. Note that for $k=0$ and $k=n$ we don’t need a parameterized Dicke state, because $k=0$ is equivalent to $|0\rangle^{\otimes n}$ and $k=n$ is the state $|1\rangle^{\otimes n}$, then the range applied to $k$ is [1,n1][1,n-1]. The curves defined by different values of $n$ and $k$ can be unified by the expression nkk(k+1)/2nk-k(k+1)/2, as depicted in the plot in the right.
XII Dicke state ansatz on noisy devices
Once a good ansatz has been identified, it is important to improve its implementation in hardware to improve the results. In reference [wang2024variational], the authors proposed a Dicke state ansatz version that is more efficient to run in the current noisy quantum hardware, reducing the number of entangling gates and also considering qubits connections constraints. Another way to improve the depth of the Dicke state circuit proposed in [wang2024variational, mukherjee2020preparing], it is the parallelization of entangling gates, with which we can optimize the circuit depth in order to create quantum circuits which are suitable for the current quantum hardware coherence time. A topic to highlight is the Dicke state sensitivity to errors of two types: bit-flip and readout errors. For instance, we can measure the impact of bit-flips, in our ansatz, modeling this type of error as described below and manipulating the probability $p$ of occurrence of a bit-flip [nielsen2010quantum]
$\rho=(1-p)\rho_{0}+p\sigma_{x}\rho_{0}\sigma_{x},$ (13)
where $\rho_{0}$ represents the density matrix related to the Dicke state ansatz. Then we can use the noisy state $\rho$ to calculate the Hamiltonian expected value through [weinberg2015lectures]
$\langle H\rangle=\textrm{tr}(\rho H),$ (14)
and with that we will be able to estimate the impact on our objective function. Both errors can destroy the main characteristics of this quantum state, that is, a constant Hamming weight, because depending of the number of bit-flips and in which qubits they will occur, this kind of error can increase or reduce the Hamming weight, which will reflect in adding Hamiltonian terms that will have a great impact on Hamiltonian expectation value. Then, it is expected that the results obtained from current noisy quantum hardware will not achieve the same level of convergence that we achieved with noiseless simulation, because bit-flips and readout errors can make parameter optimization more challenging, since noise can also induce Barren Plateaus [wang2021noise].
One last type of error that can cause issues during parameter optimization is the coherent error, more specifically, an error of the form $U(\theta+\delta\theta)$, where $\delta\theta$ represents a displacement in the original parameter. In order to mitigate the damage that the errors described above can cause, we can use some error mitigation techniques. For example, to mitigate errors in expectation values we can use ZNE [kandala2018extending, giurgica2020digital, kim2023scalable], PEC [van2023probabilistic, gupta2024probabilistic] and Probabilistic Error Amplification (PEA) [filippov2023scalable, kim2023evidence]. For readout errors, we can apply statistical corrections [nation2021scalable, van2022model], M3 technique [nation2021scalable] or Twirled Readout Error Extinction (T-REX) [van2022model]. The usage of the error mitigation techniques mentioned above, combined with advanced transpiling methods are highly recommended, in order to produce optimized quantum circuits to current hardware, improving the chances of obtaining good results. Another interesting direction in which we would like to explore CVaR as the objective function [barkoutsos2020improving, barron2024provable], once CVaR seems to be a noise-resistant loss function for variational quantum algorithms.
In order to assess the effectiveness of our simulation results on current quantum hardware, we carried out experiments utilizing the IBM QPU ibm_pittsburgh. These tests were executed in Scenario I, employing the most optimal set of parameters post-optimization. The parameters we identified ensured a $100\%$ probability of sampling the global optimum. We analyze the distribution obtained from the hardware against the simulation, which in this scenario matches the ideal result. As observed in Supplementary Figure 9, the hardware results are quite noisy and do not identify the global optima as the bit string with the highest sampling probability. Additionally, it is noteworthy that the Hamming weight of the sampled bit strings failed to maintain the Hamming established by $k=4$, since infeasible solutions appear in the results. This result aligns with the discussion in this section, where it was emphasized that in NISQ hardware, infeasible solutions may arise due to bit-flip and readout errors.
Figure 9: Scenario I: A comparison between a simulation using optimal parameters that achieves the ground state with a $100\%$ probability and the distribution obtained from hardware using the identical circuit and parameters. The states are labeled with integer numbers according to their energy levels. Thus, the ground state is labeled as 0, the first excited state as $1$, and this pattern continues similarly. In blue, we have the simulation results, while the hardware results are in red and gray. The red indicates the feasible bit strings, whereas the gray the infeasible bit strings, which do not comply with the Hamming weight constraints. Observe that the search space size is $210$, meaning that indexes greater than this indicate infeasible bit strings. The optimization level used in Qiskit Transpiler was 3.
Additionally, we calculate the expectation value of the Ising Hamiltonian for Scenario I using the Qiskit Estimator primitive over 100 executions on IBM quantum hardware. We chose the set of parameters that yielded the best outcome in simulations and incorporated these parameters into the Dicke state ansatz. Subsequently, the expectation value was calculated 100 times using the same quantum circuit. The results are shown in Supplementary Figure 10
Figure 10: Calculation of the expected value of the Hamiltonian using $100$ Qiskit Estimator executions in Scenario I. The vertical dashed black line indicates the ground state energy that we aim to reach. It is evident that the distribution is significantly distant from the optimal regime.
It is important to note that these experiments were performed using the default settings of Qiskit Sampler and Estimator, and we also considered the standard transpilation techniques adopted in Qiskit transpiler. The results presented above show what we expected to see in noisy hardware without circuit optimizations and error mitigation techniques, however these data do not contradict our simulation results, they reflect the challenge that is extract a result that is comparable with a noiseless simulation. In future research, we would like to address these problems of hardware performance with a more optimized circuit and also using the error mitigation techniques mentioned.
XIII Quantum state fidelity analysis
Figure 11: A visual analysis of the state fidelity behavior in each scenario proposed. The heatmap represents the state fidelity between of each experiment with Dicke state and CMA-ES, then each point reflects the similarity between the quantum states produced by two sets of parameters obtained at the end of the corresponding experiments. The main diagonal is always equal to 1, because it is the fidelity between the quantum states defined by the same set of parameters. The colored histogram in the Y-axis shows how the points in the heatmap are distributed in the state fidelity scale. The gray histogram represents the state fidelity distribution obtained when we compute the fidelity between the quantum states produced by the optimal set of parameters for each experiments and the reference state for each case: (a) Scenario I results. (b) Scenario II results. (c) Scenario III results.
An additional analysis that we performed was on the quantum state fidelity between states at the end of the optimization process, but also on their state fidelity with reference state. Here we only consider results obtained with Dicke state ansatz simulations using CMA-ES as optimizer, these results are summarized in Figure 15.
The state fidelity computation was done considering pure states, then we used the following mathematical expression [nielsen2010quantum]
$F(|\psi(\vec{\theta^{*}_{i}})\rangle,|\psi(\vec{\theta^{*}_{j}})\rangle)=|\langle\psi(\vec{\theta^{*}_{j}})|\psi(\vec{\theta^{*}_{i}})\rangle|^{2},$ (15)
where $\vec{\theta^{*}_{i}}$ and $\vec{\theta^{*}_{j}}$ represent two different optimal set of parameters for the states described in Table 2. To compute the state fidelity, we consider a parameterized quantum circuit that is used to compute the kernel in Quantum Support Vector Machine (QSVM) [chen2024validating], where we apply an unitary operator $U(\vec{\theta})$ that prepares the desired state and its adjoint with different set of parameters, as stated in the following expression
$|\Phi\rangle=U^{\dagger}(\vec{\theta_{j}})U(\vec{\theta_{i}})|0\rangle^{\otimes\ n}.$ (16)
Then, our estimation of fidelity will be equal to the probability of measuring $|0\rangle^{\otimes\ n}$. This fidelity estimation could also be done with the SWAP test [fanizza2020beyond].
Our finding here was that these different sets of optimal parameters lead to quantum states that have a high fidelity to the reference state. We also noted a certain level of similarity between these quantum states, as shown by the heat maps plots in the Supplementary Figure 15. However, we cannot extract useful information from the distance between the parameter sets, because it is not clear how we can define what is close or distant in the parameter space.
Perhaps, with this information in hand, we can extract some useful information about the optimization landscape to improve the success of this method. A future research line would design a neural network to provide a good initial starting set of parameters considering this information about the similarity between the states that have a high fidelity with the reference state and are similar between them.